On the first segment of my trip I drove my car 405 miles on 18 gallons of gas. On the second segment of the trip I drove 183 miles on 6 gallons of gas. What are the miles per gallon for each segment of the trip and the overall miles per gallon for the entire trip?

Answer by checkley79(3341)

405/18=22.5 mpg.
183/6=31 mpg.
(405+183)/(18+6)=588/24=24.5 mpg. ave.